# V-VAE: A Variational Auto Encoding Framework Towards Fine-Grained Control over Human-Like Chat

### v1

@@ Abstract
Qi Lin12∗, Weikai Xu1, Lisi Chen1† and Bin Dai2† 1University of Electronic Science and Technology of China 2Xaiobing.AI linqi@std.uestc.edu.cn, xuwk266@gmail.com lchen012@e.ntu.edu.sg, daibin@xiaobing.ai
11footnotetext: Work done during the internship at Xiaobing.AI.22footnotetext: Corresponding authors: Lisi Chen and Bin Dai.
@@ 1 Introduction
LLM-based chatbots are becoming increasingly popular and intelligent. Their applications are ubiquitous in a variety of specialized domains, such as online education (Kang and Kang, 2024), customer service (Park et al., 2024a), and digital human (Suo et al., 2025). These chatbots are expected to follow specific roles during human-chatbot interactions, requiring them to exhibit personalized attributes and characteristics associated with those roles (Tamoyan et al., 2024). For this purpose, existing studies aim to improve the control over persona-related attributes by employing predefined role documents through in-context learning (Chen et al., 2023), limiting the length of individual responses in the decoding phase and the number of dialogue rounds (Zhao et al., 2024), and modifying interaction styles in the training corpus (Shen, 2024). With chatbots now capable of assuming designated roles, recent efforts (Çalık and Akkuş, 2025) aim to further enhance their ability to exhibit human-like behavior in LLM-driven conversations.
@@
However, as shown in Figure 1, human-like chat imposes more stringent requirements compared to role-play chat (Tu et al., 2024) and persona-based chat (Yamashita et al., 2023), as it needs to capture more authentic human interaction histories and exhibit a broader range of conversational attributes. Specifically, human-like chat modeling poses the following three challenges: (1) Unlike static and predefined role descriptions, human-like chat unfolds dynamically, with the chatbot’s tone, lexical choices, and interaction patterns evolving for the communication process. (2) While role-play chat relies on scripted constraints and persona-based conversations use explicit labels (e.g., age, occupation) to define character traits explicitly, human-like chat involves complex and latent signals such as emotional tendencies, situational awareness, and evolving personality traits. These aspects are difficult to model from initialization and even harder to evaluate quantitatively. (3) Role-play chat exhibits a weaker dependency on such latent signals, as its generation is based on human-AI interaction for data synthesis (Kim et al., 2024) or direct teacher chatbot distillation (Hu et al., 2025). In contrast, the higher quality demands of human-like chat data, especially its fidelity to real human behavior, make it impractical to build using automated pipelines or human-machine dialogue platforms.
@@ 2.1 Human-like Chat
A variety of research has been dedicated to enhancing the human-like qualities of large language model (LLM) responses (Li et al., 2023b; Çalık and Akkuş, 2025). Techniques such as Reinforcement Learning from Human Feedback (RLHF) have significantly refined model outputs by aligning them with user preferences and expectations (Cuayáhuitl et al., 2019; Jaques et al., 2020). One prominent model, DialoGPT (Zhang et al., 2019), leverages extensive Reddit data to produce responses that closely resemble human conversation. Similarly, Meena, a multi-turn chatbot, has been optimized to achieve high dialogue coherence through metrics like Sensibleness and Specificity Average (SSA) (Durmus et al., 2023; Adiwardana et al., 2020). For benchmarks, several datasets were released to evaluate the human-likeness of chatbot in the field of human-robot interaction (Kahn Jr et al., 2007; Duan et al., 2024; Ying et al., 2025; Ou et al., 2023). Our work emphasizes the nuanced integration of real-time emotional states, relational dynamics, and interactional context between dialogue participants, dimensions often overlooked in conventional emotion-aware systems, which have been ignored in previous work.
@@
2.2 Role-play Chat
Large language models (LLMs) can demonstrate distinct behaviours and personas (Wolf et al., 2024; Andreas, 2022). Meanwhile, LLMs’ performance for role-playing and simulating human behaviour is primarily influenced by defining the persona of interest within the prompt (Aher et al., 2023; Horton, 2023; Kovač et al., 2023; Argyle et al., 2023), response length (Zhao et al., 2024) and interaction styles (Shen, 2024). For data, the methods for collecting human-chatbot conversation datasets can be categories into several tpyes, such as human-human (Durmus et al., 2023; Kim et al., 2022; Li et al., 2023a), teacher-student (Macina et al., 2023), and patient-physician (Wang et al., 2023). For downstream tasks, such as social sciences, a multitude of studies have been employing use persona prompts in LLMs to simulate human behavior (Horton, 2023; Argyle et al., 2023; Kim and Lee, 2023; Törnberg et al., 2023), while others have pointed out the lack of fidelity and diversity in such simulations (Bisbee et al., 2024; Park et al., 2024b; Wang et al., 2024; Taubenfeld et al., 2024). The LLM Roleplay framework takes a unique approach by generating diverse dialogues through persona-based interactions, further simulating human-chatbot exchanges (Takagi et al., 2025; Tamoyan et al., 2024).
@@ 3.1 Task Formulation
$-\log p_{\theta}(x|c)=-\log\sum_{z\in\mathcal{Z}}p_{\theta}(x|c,z)\cdot p_{\lambda}(z),$ (1)
@@ 3.2 Variational Auto-Encoding Mechanism
To make the objective tractable, we introduce a variational posterior distribution $q_{\phi}(z\mid x,c)$, we assume that $z$ is independent of the observable context $c$ and derive a variational upper bound of (1) for $-\log p_{\theta}(x\mid c)$:
@@ 3.4 Prior of Latent Persona Space
$\mathcal{Z}_{k}=\left\{v\mid\exists\,(x,c)\in\mathcal{D}_{\text{train}}\ (\pi_{\phi}(x,c)_{k}=v)\right\}\neq\emptyset$ (7)
where $v$ is the specific feature value of the dimension $z_{k}$ and $\mathcal{D}_{\text{train}}$ is the training corpus. However, given that the LLM interface can only reference approximately 14 conversation turns on average to infer predefined latent persona features (e.g., catchphrases, frequently-used-emojis, hobbies), feature extraction failures ($\pi_{\phi}(x,c)_{k}=\emptyset$) naturally occur due to insufficient contextual evidence. This aligns with human communication patterns — individuals do not rigidly exhibit all persona traits in every utterance, though latent traits persist beyond explicit mentions. For example, a person may frequently use a particular emoji, but not necessarily in every dialogue turn. Similarly, as illustrated in Figure 2, the AI’s hobby may not be mentioned explicitly in the conversation, yet this does not imply the absence of such a trait. Therefore, to address null values while maintaining persona consistency, we implement a probabilistic fallback mechanism(Random Sampling), where empty features are replaced by the result random sampling from the empirical prior distribution $\mathcal{Z}_{k}$ aggregated across training set. Theoretically, this resembles Markov Chain Monte Carlo (MCMC) Robert et al. (1999) initialization strategies that leverage historical distributions to guide sampling when local context lacks information.
@@ 4.1 HumanChat Data Collection
To curate a human-like chat dataset, we developed a virtual social platform where users can engage in conversations with agents created by both themselves and other users. Chat sessions are logged and stored on the server***The platform will be de-anonymized, and the dataset will be released alongside the official publication of this paper.. Prior to dataset construction, we apply a preprocessing pipeline to filter out sensitive information. This includes but is not limited to personal details such as names, phone numbers, and locations. For highly active agents, we intentionally subsample their conversations to mitigate redundancy in the dataset. Using the cleaned sessions, we employ 30 annotators from diverse backgrounds to identify and refine low-quality responses—defined as utterances that deviate from typical human conversational patterns. Annotators rewrite these turns to align with natural human communication, with the rewritten response designated as the target $x$ and the preceding context as $c$. While rewriting, annotators inherently infuse their individual styles into the responses; these stylistic elements, though unrepresented in $c$, serve as the latent variable $z$ that our encoder aims to recover. The final dataset is divided into two subsets: HumanChatData for training and HumanChatBench for evaluation. Detailed statistics of the dataset are presented in Table 1.
@@ 4.2 Experiment Setup
Benchmarks. Our method is evaluated on HumanChatBench and DialogBench. DialogBench Ou et al. (2023), a multi-task benchmark for evaluating dialogue systems, comprises 12 tasks assessing LLMs’ abilities to exhibit human-like conversational behaviours. We adopted a subset of seven tasks based on their semantic relevance and operational fidelity to our experimental framework. While HumanChatBench emphasizes fine-grained linguistic behaviors to assess human-likeness, DialogBench focuses on higher-level capabilities such as knowledge comprehension and offense detection, which are more comprehensive but less average dialogue turns compared with HumanChatBench (7.58 vs. 14.4).
@@
Baselines. We conduct experiments to evaluate the performance of different methods based on the following settings: (1)+FT, standard finetuning using our originally collected and annotated dataset; (2)+P+FT, context-aware augmentation that incorporates latent space derived from contextual utterances into the training data; and (3)+SP+FT an enhanced variant of the second approach employing post-sampling refinement procedures specifically for scenarios where latent spaces exhibit null values. The first method establishes baseline performance through conventional supervised learning. The second approach enhances model adaptability by injecting discourse-specific personality signatures obtained via Verbal VAE encoding of dialogue contexts. The third methodology addresses sparse latent space conditions through sampling of persona-absent instances from the prior of latent space, thereby improving robustness against incomplete persona manifestations. We adopt LLaMA3-8B (Grattafiori et al., 2024), Qwen-7B (Bai et al., 2023), and Qwen-14B as strong open-source LLM baselines, and include Doubao, GPT-4o-mini (Hurst et al., 2024), and StepAI as commercial API-based models to provide a comprehensive comparison across different settings.
@@ 4.4 Ablation Study
We perform the following ablation tests to validate the effect of each component of the persona structure: (1) Remove the talking style information about the chatbot(-talking); (2) Remove the interaction style between the chatbot and the user(-interaction). (3) Remove some information about personal style (-personal). We conducted LoRA fine-tuning experiments using three versions of data on two models, LLaMA3-8B and Qwen-7B, and evaluated them based on the Validation Loss and HumanChatBench metrics. The results are shown in Table LABEL:tab:ablation. We observe that: (1) the components in our persona dataset exhibit a descending order of importance: talking > interaction > personal. Removing more critical components leads to higher validation loss, with experimental results showing progressively lower losses when removing talking (1.73), interaction (1.71), and personal (1.69) data respectively. (2) On HumanChatBench, the talking style space has a greater impact on the fine-grained control of response language(e.g., its removal leads to larger deviations from the target on the CP metric). In contrast, on the EC metric, models perform better without talking style than with, suggesting that random sampling for null values may introduce more noise in this dimension compared to others.
@@ 4.5 Further Discussion
Compare with close-source LLMs. We further evaluated close-source LLMs under both zero-shot and few-shot settings, with the comprehensive results documented in Table 4. In the zero-shot configuration, the close-source LLMs exhibit strong metric on HumanChatBench indices (e.g., GPT-4o-mini performs 84.08 on CP and 88.66 on EC) owing to their robust instruction-following capabilities. However, we posit that optimal performance should align with human-annotated ground truth metrics (i.e., proximity to the Target reference values). Under few-shot conditions, the API demonstrates adaptive behaviour by selectively adhering to provided persona features rather than rigidly enforcing all characteristics, resulting in improved alignment with Target (e.g., GPT-4o-mini shows a reduction of 66.28 on CP of Target deviation under 2-shot learning compared to the zero-shot scenario.). Nevertheless, its performance remains statistically inferior to our fine-tuned model across all HumanChatBench metrics. This systematic comparison highlights two critical insights: 1) Instruction fidelity does not guarantee appropriate human-like persona consistency, rather, strong instruction-following tendencies may lead chatbots more AI-like. 2) Compared to prompt engineering, parameter optimization offers superior control over nuanced persona adaptation, especially in capturing subtle and context-dependent identity cues.
Performance on structured persona and unstructured persona. To validate the rationality of the persona’s structural design, we first employed Doubao’s self-diagnostic capability to analyze: (1) the underlying causes, and (2) the AI’s distinctive characteristics that generated the annotated utterance without structural constraints. We then performed comparative LoRA fine-tuning experiments across two models using this dataset, with comprehensive evaluations conducted. As is shown in Table LABEL:tab:unstructured, our experimental analysis reveals a great difference between structured and non-structured data settings across both models. For the two models, non-structured data achieved substantially lower validation losses (0.54 and 0.63, respectively). However, structured data configurations exhibited closer alignment with the golden metrics for persona consistency: CP (8.23–9.65 vs. 43.86-50.81 with Target 8.57) and HM (1.15–1.48 vs. 3.98-9.31 with Target 2.16). The case study reveals that the unstructured persona predominantly focuses on the immediate discourse context—specifically analyzing why the AI produces a given utterance within its logical framework, rather than attributing responses to the AI’s characteristics. As detailed in the Appendix, this approach prioritizes contextual reasoning over persona-driven trait associations.

### v2

@@ Abstract
Qi Lin1, Weikai Xu1, Lisi Chen1∗ and Bin Dai2∗ 1University of Electronic Science and Technology of China 2Qingyao Intelligence linqi@std.uestc.edu.cn, xuwk266@gmail.com lchen012@e.ntu.edu.sg, daibin@jinyao.tech
11footnotetext: Corresponding authors: Lisi Chen and Bin Dai.
@@ 1 Introduction
LLM-based chatbots are becoming increasingly popular and intelligent. Their applications are ubiquitous in a variety of specialized domains, such as online education (chang2025review), customer service (park2024align), and digital human (suo2025digital). These chatbots are expected to follow specific roles during human-chatbot interactions, requiring them to exhibit personalized attributes and characteristics associated with those roles (tamoyan2024llm). For this purpose, existing studies aim to improve the control over persona-related attributes by employing predefined role documents through in-context learning (chen2023places), limiting the length of individual responses in the decoding phase and the number of dialogue rounds (zhao2024long), and modifying interaction styles in the training corpus (shen2024rethinking). With chatbots now capable of assuming designated roles, recent efforts (ccalik2025enhancing) aim to further enhance their ability to exhibit human-like behavior in LLM-driven conversations.
@@
However, as shown in Figure 1, human-like chat imposes more stringent requirements compared to role-play chat (tu2024charactereval) and persona-based chat (yamashita2023realpersonachat), as it needs to capture more authentic human interaction histories and exhibit a broader range of conversational attributes. Specifically, human-like chat modeling poses the following three challenges: (1) Unlike static and predefined role descriptions, human-like chat unfolds dynamically, with the chatbot’s tone, lexical choices, and interaction patterns evolving for the communication process. (2) While role-play chat relies on scripted constraints and persona-based conversations use explicit labels (e.g., age, occupation) to define character traits explicitly, human-like chat involves complex and latent signals such as emotional tendencies, situational awareness, and evolving personality traits. These aspects are difficult to model from initialization and even harder to evaluate quantitatively. (3) Role-play chat exhibits a weaker dependency on such latent signals, as its generation is based on human-AI interaction for data synthesis (kim2024understanding) or direct teacher chatbot distillation (hu2025efficient). In contrast, the higher quality demands of human-like chat data, especially its fidelity to real human behavior, make it impractical to build using automated pipelines or human-machine dialogue platforms.
@@ 2.1 Human-like Chat
A variety of research has been dedicated to enhancing the human-like qualities of large language model (LLM) responses (li2023reinforcement; ccalik2025enhancing). Techniques such as Reinforcement Learning from Human Feedback (RLHF) have significantly refined model outputs by aligning them with user preferences and expectations (cuayahuitl2019deep; jaques2020human). One prominent model, DialoGPT (zhang2019dialogpt), leverages extensive Reddit data to produce responses that closely resemble human conversation. Similarly, Meena, a multi-turn chatbot, has been optimized to achieve high dialogue coherence through metrics like Sensibleness and Specificity Average (SSA) (durmus2023towards; adiwardana2020towards). For benchmarks, several datasets were released to evaluate the human-likeness of chatbot in the field of human-robot interaction (kahn2007human; duan2024hlb; ying2025benchmarking; ou2023dialogbench). Our work emphasizes the nuanced integration of real-time emotional states, relational dynamics, and interactional context between dialogue participants, dimensions often overlooked in conventional emotion-aware systems, which have been ignored in previous work.
@@
2.2 Human-likeness
Prior research (danesi2017language; le2017evolution) has established foundational sociolinguistic influences spanning identity construction, digital communication patterns, and multimodal semiotics like emoji usage. While large language models (LLMs) derive training data predominantly from naturalistic texts, their inherent one-to-many deployment paradigm—wherein a single model serves heterogeneous user bases—engenders systematic objectivity in politeness conventions and generates perceptibly AI-like linguistic outputs. Consequently, significant scholarly investigations have quantified sociolinguistic divergences between LLM and human communication(schneider2024sociolinguist). These variations manifest across three core dimensions: Talking style, encompassing emoji interpretation and deployment dynamics(zheng2025irony) and measurable disparities in grammatical and rhetorical composition (schneider2024sociolinguist; reinhart2025llms); Interaction patterns, characterized by relational alignment failures during longitudinal interactions(altenburger2024examining); Personal attributes, evidenced in absent human-like psycholinguistic properties in extended discourse(seals2023long). Our feature space design is therefore grounded in these empirically validated dimensions of cross-modal variation, systematically addressing each categorical divergence through targeted representation learning.
@@ 3.1 Task Formulation
$\log p_{\theta}(x|c)=-\log\sum_{z\in\mathcal{Z}}p_{\theta}(x|c,z)\cdot p_{\lambda}(z),$ (1)
@@ 3.2 Variational Auto-Encoding Mechanism
To make the objective tractable, we introduce a variational posterior distribution $q_{\phi}(z\mid x,c)$, we assume that $z$ is independent of the observable context $c$ and derive a variational upper bound of (1) for $\log p_{\theta}(x\mid c)$:
@@ 3.4 Prior of Latent Persona Space
$\mathcal{Z}_{k}=\left\{v\bigm|\exists(x,c)\in\mathcal{D}_{\text{train}},\ \pi_{\phi}(x,c)_{k}=v\neq\emptyset\right\}$ (7)
where $v$ is the specific feature value of the dimension $z_{k}$ and $\mathcal{D}_{\text{train}}$ is the training corpus. However, given that the LLM interface can only reference approximately 14 conversation turns on average to infer predefined latent persona features (e.g., catchphrases, frequently-used-emojis, hobbies), feature extraction failures ($\pi_{\phi}(x,c)_{k}=\emptyset$) naturally occur due to insufficient contextual evidence. This aligns with human communication patterns — individuals do not rigidly exhibit all persona traits in every utterance, though latent traits persist beyond explicit mentions. For example, a person may frequently use a particular emoji, but not necessarily in every dialogue turn. Similarly, as illustrated in Figure 2, the AI’s hobby may not be mentioned explicitly in the conversation, yet this does not imply the absence of such a trait. Therefore, to address null values while maintaining persona consistency, we implement a probabilistic fallback mechanism (Random Sampling), where empty features are replaced by the result random sampling from the empirical prior distribution $\mathcal{Z}_{k}$ aggregated across training set. Theoretically, this resembles Markov Chain Monte Carlo (MCMC) robert1999monte initialization strategies that leverage historical distributions to guide sampling when local context lacks information.
@@ 4.1 HumanChat Data Collection
To curate a human-like chat dataset, we developed a virtual social platform where users can engage in conversations with agents created by both themselves and other users. Chat sessions are logged and stored on the server, the platform has been de-anonymized, and the HumanChatData***https://huggingface.co/datasets/me-no-money/HumanChatData could be publicly accessable. Prior to dataset construction, we apply a preprocessing pipeline to filter out sensitive information. This includes but is not limited to personal details such as names, phone numbers, and locations. For highly active agents, we intentionally subsample their conversations to mitigate redundancy in the dataset. Using the cleaned sessions, we employ 30 annotators(details of the anatators’ demographic can be found in appendix) from diverse backgrounds to identify and refine low-quality responses—defined as utterances that deviate from typical human conversational patterns. Annotators rewrite these turns to align with natural human communication, with the rewritten response designated as the target $x$ and the preceding context as $c$. While rewriting, annotators inherently infuse their individual styles into the responses; these stylistic elements, though unrepresented in $c$, serve as the latent variable $z$ that our encoder aims to recover. The final dataset is divided into two subsets: HumanChatData for training and HumanChatBench for evaluation. Detailed statistics of the dataset are presented in Table 1.
@@ 4.2 Experiment Setup
Benchmarks. Our method is evaluated on HumanChatBench and DialogBench. DialogBench ou2023dialogbench, a multi-task benchmark for evaluating dialogue systems, comprises 12 tasks assessing LLMs’ abilities to exhibit human-like conversational behaviours. We adopted a subset of seven tasks based on their semantic relevance and operational fidelity to our experimental framework. While HumanChatBench emphasizes fine-grained linguistic behaviors to assess human-likeness, DialogBench focuses on higher-level capabilities such as knowledge comprehension and offense detection, which are more comprehensive but less average dialogue turns compared with HumanChatBench (7.58 vs. 14.4).
@@
Baselines. We conduct experiments to evaluate the performance of different methods based on the following settings: (1)+FT, standard finetuning using our originally collected and annotated dataset; (2)+P+FT, context-aware augmentation that incorporates latent space derived from contextual utterances into the training data; and (3)+SP+FT an enhanced variant of the second approach employing post-sampling refinement procedures specifically for scenarios where latent spaces exhibit null values. The first method establishes baseline performance through conventional supervised learning. The second approach enhances model adaptability by injecting discourse-specific personality signatures obtained via Verbal VAE encoding of dialogue contexts. The third methodology addresses sparse latent space conditions through sampling of persona-absent instances from the prior of latent space, thereby improving robustness against incomplete persona manifestations. We adopt LLaMA3-8B (grattafiori2024llama), Qwen-7B (bai2023qwen), and Qwen-14B as strong open-source LLM baselines, and include Doubao, GPT-4o-mini (hurst2024gpt), and StepAI as commercial API-based models to provide a comprehensive comparison across different settings.
@@ 4.3 Main Results
Model
Validation
Loss $\downarrow$
HumanChatBench (%)
CP EC HM
LLaMA3-8B
+SP+FT 1.69 8.23 2.83 1.15
-talking 1.73 16.73 7.09 1.08
-interaction 1.71 9.72 3.64 1.48
-personal 1.69 10.12 2.83 1.55
Qwen-7B
+SP+FT 1.87 9.65 2.90 1.48
-talking 1.94 19.57 3.85 1.21
-interaction 1.92 9.11 3.10 1.42
-personal 1.90 10.05 3.64 1.42
Target - 8.57 6.41 2.16
Table 3: Ablation study across different components of the structured sampled-persona.
@@ 4.4 Ablation Study
We perform the following ablation tests to validate the effect of each component of the persona structure: (1) Remove the talking style information about the chatbot(-talking); (2) Remove the interaction style between the chatbot and the user(-interaction). (3) Remove some information about personal style (-personal). We conducted LoRA fine-tuning experiments using three versions of data on two models, LLaMA3-8B and Qwen-7B, and evaluated them based on the Validation Loss and HumanChatBench metrics. The results are shown in Table 3. We observe that: (1) the components in our persona dataset exhibit a descending order of importance: talking > interaction > personal. Removing more critical components leads to higher validation loss, with experimental results showing progressively lower losses when removing talking (1.73), interaction (1.71), and personal (1.69) data respectively. (2) On HumanChatBench, the talking style space has a greater impact on the fine-grained control of response language(e.g., its removal leads to larger deviations from the target on the CP metric). In contrast, on the EC metric, models perform better without talking style than with, suggesting that random sampling for null values may introduce more noise in this dimension compared to others.
@@ 4.5 Further Discussion
Compare with close-source LLMs. We further evaluated close-source LLMs under both zero-shot and few-shot settings, with the comprehensive results documented in Table 4. In the zero-shot configuration, the close-source LLMs exhibit strong metric on HumanChatBench indices, owing to their robust instruction-following capabilities. However, we posit that optimal performance should align with human-annotated ground truth metrics (i.e., proximity to the Target reference values). Under few-shot conditions, the API demonstrates adaptive behaviour by selectively adhering to provided persona features rather than rigidly enforcing all characteristics, resulting in improved alignment with Target. Nevertheless, its performance remains statistically inferior to our fine-tuned model across all HumanChatBench metrics. This systematic comparison highlights two critical insights: 1) Instruction fidelity does not guarantee appropriate human-like persona consistency, rather, strong instruction-following tendencies may lead chatbots more AI-like. 2) Compared to prompt engineering, parameter optimization offers superior control over nuanced persona adaptation, especially in capturing subtle and context-dependent identity cues.
Model
Validation
Loss $\downarrow$
HumanChatBench (%)
CP EC HM
LLaMA3-8B
+SP+FT 1.69 8.23 2.83 1.15
+unstructured 0.54 43.86 49.39 3.98
Qwen-7B
+SP+FT 1.87 9.65 2.90 1.48
+unstructured 0.63 50.81 35.29 9.31
Target 8.57 6.41 2.16
Table 5: Performance comparison across structured persona and unstructured persona.
Performance on structured persona and unstructured persona. To validate the rationality of the persona’s structural design, we first employed Doubao’s self-diagnostic capability to analyze: (1) the underlying causes, and (2) the AI’s distinctive characteristics that generated the annotated utterance without structural constraints. We then performed comparative LoRA fine-tuning experiments across two models using this dataset, with comprehensive evaluations conducted. As is shown in Table 5, our experimental analysis reveals a great difference between structured and non-structured data settings across both models. For the two models, non-structured data achieved substantially lower validation losses (0.54 and 0.63, respectively). However, structured data configurations exhibited closer alignment with the golden metrics for persona consistency: CP (8.23–9.65 vs. 43.86-50.81 with Target 8.57) and HM (1.15–1.48 vs. 3.98-9.31 with Target 2.16). The case study reveals that the unstructured persona predominantly focuses on the immediate discourse context—specifically analyzing why the AI produces a given utterance within its logical framework, rather than attributing responses to the AI’s characteristics. As detailed in the Appendix, this approach prioritizes contextual reasoning over persona-driven trait associations.